# MMGT: Motion Mask Guided Two-Stage Network for Co-Speech Gesture Video Generation

### v1

@@ IV-A Experimental setup
Evaluation Metrics. The Fréchet Gesture Distance (FGD) [49] is employed to measure the distribution gap between real and generated gestures in the feature space, while Diversity (Div.) [50] quantifies the variability of the generated gestures. Both metrics are computed using an auto-encoder trained on PATS pose data. For evaluating video quality, we leverage Fréchet Video Distance (FVD)[51] and Fréchet Image Distance (FID) with the I3D classifier [52] pre-trained on Kinetics-400 [52]. Furthermore, PSNR[53], SSIM[54] and LPIPS are used to assess pixel-level similarity, ensuring that the generated video closely resembles the ground truth. We also use the SyncNet[55] to calculate Sync-C and Sync-D for valid the audio-lip synchronization accuracy.
Implementation Details. We train our overall framework on four speakers jointly in two stages. 1) For the SMGA-Network: We collect keypoints and expand them into motion features $x^{i}_{0}\in\mathbb{R}^{402}$ for each frame. For audio feature extraction, we adopt the feature extraction strategy outlined in [11], we set the batch size to 256 and utilize the Adan optimizer with a learning rate of $2\times 10^{-4}$, a weight decay of 0.02, and 3,400 training iterations. The maximum sampling step $T$ is set to 50, and the loss weights In Eq. 18, are configured as $\lambda_{\text{f}}=3$ and $\lambda_{\text{g}}=1$. 2) The multimodal video generation model is trained in two step. The first step comprises 29,580 steps with a batch size of 2, while the second step consists of 32,500 steps with a batch size of 1. Both step share consistent parameter settings, including a video resolution of $512\times 512$. To enhance the quality and robustness of video generation, all input modalities are randomly removed with a probability of 5% during the training process. The learning rate for both stages is set to $1\times 10^{-5}$, and optimization is performed using the Adam optimizer.[56].
Experiments encompassing both training and inference were carried out on a computational platform featuring 8 NVIDIA A6000 GPUs. Our training strategy is shown in Algorithm 2. As shown in the Fig.1, in order to make a fair comparison during the inference process, our framework only uses audio and inference images as inputs to the model without introducing other a priori features.
TABLE I: Quantitative comparison of speech-driven gesture video generation methods in terms of gesture, face, lip and overall video quality metrics. This table compares the performance of our proposed method with baseline models (MM-Diffusion[2], EDGE + Animate-anyone [48, 47], EDGE + hallo[48, 21] in audio-driven and pose-driven video generation tasks.
@@
MM-Diffusion[2] 141.76 0.0998 39.28 693.311 63.150 27.94 0.1312 0.6112 13.75 0.82 86.18 1007.01 27.93 0.1667 0.7058
EDGE.+Animate-anyone[48, 47]. 6.72 0.0855 4.20 176.915 17.36 29.76 0.4188 0.2670 12.27 2.01 6.91 410.92 31.16 0.6152 0.2374
EDGE.+Hallo[48, 21]. 7.57 0.0846 5.72 245.095 24.172 29.54 0.4158 0.2695 12.09 2.28 8.66 398.10 30.44 0.6056 0.2405
@@
TABLE II: Quantitative Comparison of Video-Driven Gesture Video Generation Methods in Terms of Gesture, Face, Lip, and Overall Video Quality Metrics. This table compares the performance of our proposed method with baseline models Animate-anyone [47] in pose video-driven video generation tasks.
@@ IV-B Quantitative result
To enhance evaluation efficiency, we standardized the resolution of all videos to $256\times 256$ for assessment. A comprehensive evaluation of various metrics was conducted using a consistent test set to ensure fair and reliable comparisons.
Audio Driver. Currently, open-source training code for audio-driven speech gesture video generation methods is relatively limited, and the training code for some methods, such as [8], [22], [57], [1], and [58], is not publicly available, preventing direct comparisons on the same dataset. To evaluate the advantages of our approach, we modify and integrate some of the existing algorithms, adapting them to the task at hand, and perform uniform training and inference using the same training and test sets.
We constructed two benchmark models of two-stage speech-driven upper body speaking based on current state-of-the-art algorithms such as EDGE[48], Animate-anyone[47] and Hallo [21]. In making modifications to EDGE[48], we refer to S2G-MDD[11] scheme to adapt it to the current dataset. First, we use DWPose [41] to extract the keypoint sequences on the PATS dataset [27, 28, 29] and construct a paired dataset of gesture sequences and audio as training data for EDGE[48]. For Hallo[21], we add a pose-driven pose guider [47] to it so that it can introduce pose videos during training. In terms of mask acquisition, since Hallo[21] can only acquire static masks for faces, and faces in speech videos are small and cannot acquire masks efficiently, we adjusted the mask acquisition method by using YOLOv5[59] to acquire static masks for different regions of the human body, making it more suitable for the task of generating videos of the upper half of the body.
In the second phase, we train Animate-anyone[47] using pose videos and GT videos, while we train Hallo[21] using pose videos, GT videos, and audio. In terms of inference, we use the same test set, first generating videos of DWPose keypoints [41] using EDGE, and then using these keypoints for each of Hallo [21] and Animate-anyone [47] video rendering respectively.
As shown in Table I, our method significantly outperforms other approaches in overall video quality. Specifically, our FVD score improves from 398.10 (for EDGE+Hallo) to 230.89, highlighting a notable enhancement. Moreover, our model achieves the best results in key metrics such as PSNR, SSIM, and LPIPS. In lip synchronization, our method excels with superior performance in both Syn-C and Syn-D metrics. For facial quality, our approach also delivers the best results. In terms of gesture generation, we achieve a clear improvement in the FGD metric. Although the Div metric may not accurately reflect pose generation quality when significant jitter is present, it is less indicative in this context. This is further demonstrated in the qualitative experiments.
Pose video-driven generation. We adapted the approach by replacing the audio-driven pose videos with whole-body keypoint sequences extracted from the PATS dataset[27, 28, 29] using DWPose [41] during the inference process for the video-driven gesture generation task. As shown in Table II, our method outperforms the current state-of-the-art method, Animate-anyone[47], across multiple evaluation metrics, including FID, FVD, PSNR, SSIM, and LPIPS, for both the overall video quality and the facial region. Specifically, our model shows improvements in FID, FVD, and LPIPS, achieving better perceptual similarity and video coherence. Furthermore, the visual results presented in Fig. 7 demonstrate that our approach, MMGT, excels in maintaining structural integrity and identity consistency, especially in localized regions like hands.
@@
TABLE III: The quantitative results of the ablation study highlight the impact of the key modules on gesture generation, video quality and audiovisual synchronization metrics.
@@
Qualitative Comparative Experiment. We show a qualitative comparison of the upper body speech-driven video generation model on the PATS dataset, as seen in Fig. 6, where EDGE + Animate-anyone [48, 47] hand movements are delayed and unnatural, and lip synchronization is not aligned. edge + hallo[48, 21] face There is noticeable distortion, hand distortion, and a lot of artifacts in the background. In contrast, our proposed method generates videos that are closer to GT, with more natural and synchronized gestures, accurate lip alignment, and no artifacts in the background.
We also demonstrate a qualitative comparison of video generation models driven by upper body pose videos on the PATS dataset, as shown in Fig.7, where Animate-anyone[47] generated video fails to maintain consistency with ground truth in the hand region. In contrast, our approach yields more accurate and visually consistent results, effectively maintaining the integrity of facial and hand movements.
@@
Lips Synchronization with Audio. For the qualitative comparison of lip synchronization, Fig. 8 illustrates the lip synchronization results for the word ’negative’ in the audio phrase. Our method (MMGT) produces lip shapes that are closely aligned with the Ground Truth, accurately capturing both the timing and subtle dynamics of the word. In contrast, EDGE + Animate-anyone [48, 47] and MM-Diffusion[2] show noticeable mismatches in lip movements, resulting in poor synchronization with the audio. EDGE + Hallo [48, 21] exhibits large distortions and artifacts in the lips. Our ablation model (MMGT w/o SMGA), the generated lip shape is nearly stationary and significantly deviates from the ground truth. Additionally, we provide quantitative metrics for lip synchronization in Table I and Table III, which include Syn-C and Syn-D to evaluate synchronization accuracy and diversity, respectively. The results further validate the robustness and superiority of our method over the baseline and ablation variants.
@@
Figure 10: Presentation and qualitative comparison of overall head diversity maps based on the PATS dataset
Figure 11: Hands Detailed Comparisons. Detailed comparisons of hand motion among MM-Diffusion[2], EDGE + Animate-anyone [48, 47], EDGE + Hallo[48, 21], our MMGT model, and our ablation model (MMGT w/o SMGA).
@@
Effectiveness of Motion Masks in Video Generation. Fig. 9 illustrates the qualitative comparison of video generation results under the same input conditions using still masks and motion masks. Two representative frames are selected to showcase the effectiveness of the proposed method. As shown, the incorporation of motion masks leads to significant improvements in the generated video quality, particularly in the facial and lip regions, where finer details and movement synchronization are better captured. By leveraging the dynamic information provided by motion masks, the model generates more realistic and coherent videos compared to using still masks. To provide a comprehensive assessment, we further validate these observations through ablation experiments using quantitative metrics, as shown in Table III.
@@
Evaluation of Facial Motion Diversity and Synchronization. We conducted a detailed comparison of facial motion diversity and synchronization for individual test videos, evaluating Ground Truth, EDGE + Animate-anyone [48, 47], EDGE + Hallo [48, 21], our method MMGT, and the ablation version of our method (MMGT w/o SMGA and MMGT w/o $\mathcal{L}_{f}$). As shown in Fig. 10, the motion diversity heatmap (DIV.IMG) on the left side visualizes the magnitude of the motions, where blue represents the smallest motions, green and yellow indicate moderate motions, and red highlights the most significant motions. From the comparison of the diversity maps, our method captures the most pronounced and accurate lip movements that closely match the ground truth.
The qualitative comparison on the right further highlights the differences between the various methods. Our approach not only generates accurate and well-synchronized lip shapes, but also preserves subtle micro-expressions, resulting in output that closely matches the Ground Truth. In contrast, the variant without SMGA exhibits almost no lip movement, underscoring the critical role of the SMGA module in achieving precise lip synchronization with the audio. Similarly, the ablated model w/o $\mathcal{L}_{f}$, which lacks the facial loss function, produces rigid face and less distinct lip motion, underscoring the importance of the facial loss for natural expression generation. EDGE + Animate-anyone [48, 47] and EDGE + Hallo [48, 21] struggle to generate synchronized lips with detailed facial movements. In comparison, our method excels in both motion diversity and quality. These results clearly validate the effectiveness of our full model design in generating realistic, synchronized upper-body videos.
@@
Detailed Comparison of Hand Motion Across Methods. We highlight the performance of different methods in reproducing hand motion. As shown in Fig. 11, MM-Diffusion [2] generates noticeable artifacts and blurred hand shapes, failing to capture the fine details and consistency of hand movements. Similarly, EDGE + Animate-anyone [48, 47] and EDGE + Hallo [48, 21] struggle with accurate hand positioning and shape generation, resulting in significant discrepancies from the Ground Truth. In contrast, our method demonstrates superior fidelity, accurately capturing realistic hand gestures with precise motion and clear structural details, closely aligning with the Ground Truth. This comparison emphasizes the effectiveness of our method in generating coherent and visually accurate hand motions in video synthesis.
@@
We conducted ablation studies on the PATS dataset to evaluate the influence of key components in our framework, with the quantitative results presented in Table III. In this table, ”w/o SMGA” indicates the replacement of our SMGA network with a baseline keypoint generation method (EDGE [48]), ”w/o ($\mathcal{L}_{f}$)” corresponds to removing the facial loss ($\mathcal{L}_{f}$) in the first stage, ”w/o Motion Mask” signifies the omission of the motion mask in the MM-HAA module, and “w/o Audio” replaces audio cross-attention with self-attention to negate the effect of audio inputs.
We replaced the dual-motion block in the SMGA network with a single-motion block while using $x_{t}$ as input to the ablation network in the first stage, which resulted in a significant decrease in lip synchronization and overall motion quality. As shown in Table III, the Syn-C metric increased significantly from 10.16 to 11.83, indicating a decrease in the credibility of lip synchronization, and the Syn-D metric increased from 4.51 to 5.38, reflecting a weakening of the temporal consistency of lip movements. Gesture and facial quality also decreased, as evidenced by higher FGD and FVD scores, highlighting the importance of SMGA in maintaining robust lip synchronization (Fig.11 and Fig.10).
@@
Similarly, substituting the motion-specific facial loss $\mathcal{L}_{f}$ with a generalized loss leads to a pronounced decrease in lip synchronization accuracy. The Syn-C metric rises from 10.16 to 11.11, and the Syn-D metric increases from 4.51 to 4.65, signifying diminished articulation precision and temporal coherence between the lips and audio. As shown in Fig. 10, this degradation visibly manifests as weaker lip movements aligned with the spoken content.
Eliminating the motion mask in the MM-HAA module further degrades performance. The facial FVD score increases to 205.08, while the overall video FVD soars to 378.32, indicating deteriorating temporal coherence and video fidelity. Moreover, metrics related to facial detail (PSNR, KVD, and LPIPS) also show declines, indicating loss of detail and perceptual quality. Although lip synchronization remains relatively stable—thanks to shared pose-driven inference—the enhancements in other metrics reaffirm the value of dynamic masking for preserving both global video quality and fine-grained details.
@@
Removing the audio input entirely and relying solely on keypoint-driven generation leads to substantial performance drops in both facial and overall metrics (e.g., FVD and FID), emphasizing the essential role of audio in enhancing visual quality and temporal consistency. Furthermore, the notable decrease in lip synchronization highlights the necessity of audio-based guidance for achieving realistic, accurate lip movements.
@@
When a static mask replaces the motion mask in the MM-HAA module, face-specific metrics such as SSIM and PSNR decrease (from 0.3945 to 0.3775 and from 29.47 to 29.42, respectively), while the facial FVD score increases from 99.38 to 267.10, indicating a marked loss of detail in the face region. Gesture coherence also degrades (FGD rises from 7.04 to 7.089), and the overall video FVD increases from 230.89 to 286.389, pointing to diminished temporal consistency and video quality (Fig. 9).
Collectively, these ablation studies highlight the crucial roles played by the SMGA network, the motion-specific facial loss $\mathcal{L}_{f}$, the motion mask, and audio-based guidance. Each component significantly contributes to generating high-quality, semantically rich, and temporally coherent facial and gesture videos, as demonstrated by the superior performance of the MMGT network.

### v2

@@ IV-A Experimental Setup
Evaluation Metrics. The Fréchet Gesture Distance (FGD) [57] is used to measure the distribution gap between real and generated gestures in the feature space, while Diversity (Div.) [29] quantifies the variability of the generated gestures. Both metrics are calculated using an auto-encoder trained on PATS pose data. For assessing video quality, we leverage Fréchet Video Distance (FVD)[48], kernel video distance (KVD), and Fréchet Image Distance (FID) with the I3D classifier [8] pre-trained on Kinetics-400 [8]. Additionally, PSNR[19], SSIM[52], and LPIPS are employed to evaluate pixel-level similarity, ensuring that the generated video closely matches the ground truth. We also use SyncNet[13] to compute Sync-C and Sync-D to validate audio-lip synchronization accuracy.
Implementation Details. We train our overall framework on four speakers jointly in two stages. (1) For the SMGA-Network: We collect keypoints and expand them into pose sequence $x^{i}_{0}\in\mathbb{R}^{402}$ for each frame. For audio feature extraction, we adopt the feature extraction strategy outlined in [17], we set the batch size to 256, and use the Adam optimizer with a learning rate of $2\times 10^{-4}$, a weight decay of 0.02, and 3,400 training epoch. The sampling step $T$ is set to 50.
(2) The multimodal video generation model is trained in two training processes. The first process involves 29,580 steps with a batch size of 2, while the second process includes 32,500 steps with a batch size of 1. Both processes have consistent parameter settings, including a video resolution of $512\times 512$. To improve the quality and robustness of video generation, all input modalities are randomly removed with a probability of 5% during training. The learning rate for both stages is set to $1\times 10^{-5}$, and optimization is done using the Adam optimizer.[22]. All training experiments were conducted on a computing platform equipped with eight NVIDIA A6000 GPUs, whereas inference was carried out on a single NVIDIA A6000 GPU. The complete training procedure is summarized in Algorithm 2. In addition, we report a comprehensive comparison of the computational overhead among our method and several representative baselines, including Animate-Anyone[20], MimicMo[60], and Hallo[54]. The results are summarized in Table I.
TABLE II: Quantitative comparison of co-speech gesture video generation methods in terms of gesture, face, lip, and overall video quality metrics. This table compares the performance of our proposed method with baseline models ($M$-Diffusion[43], EDGE + Animate-anyone [47, 20], EDGE + Hallo[47, 54], and DiffGest + MimicMo[63, 60] in audio-driven and pose-driven video generation tasks. The bold text indicates the optimal results, and the underlined text indicates results that were obtained.
@@
$M$-Diffusion[43] 141.76 0.0998 39.28 693.31 63.150 27.94 0.1312 0.6112 13.75 0.82 86.18 1007.01 27.93 0.1667 0.7058
EDGE + Animate-anyone[47, 20] 6.72 0.0855 4.20 176.92 17.36 29.76 0.4188 0.2670 12.27 2.01 6.91 410.92 31.16 0.6152 0.2374
EDGE + Hallo[47, 54] 7.57 0.0846 5.72 245.10 24.17 29.54 0.4158 0.2695 12.09 2.28 8.66 398.10 30.44 0.6056 0.2405
DiffGest + MimicMo[63, 60] 13.17 0.0854 5.78 239.53 18.37 29.68 0.4121 0.2837 12.48 1.21 9.40 246.63 31.13 0.6408 0.2390
@@
TABLE III: Quantitative Comparison of Pose-Driven Gesture Video Generation Methods in Terms of Gesture, Face, Lip, and Overall Video Quality Metrics. This table compares the performance of our proposed method with baseline models Animate-anyone [20] and MimicMo[60] in pose video-driven video generation tasks.
@@
TABLE IV: Sensitivity analyses concerning the loss weights of the first stage audio-to-pose model.
$\lambda_{F}:\lambda_{B}$ 1:1 1:2 1:3 1:4
FGD ($\downarrow$) 6.705 7.296 7.268 7.369
Syn-D($\uparrow$) 4.308 4.268 4.409 4.361
@@ IV-B Quantitative Result
To improve evaluation efficiency, we standardized the resolution of all videos to $256\times 256$ for assessment. A comprehensive evaluation of different metrics was conducted using a consistent test set to ensure fair and reliable comparisons. Currently, open-source training code for co-speech gesture video generation methods is relatively limited, and the training code for some methods, such as [18], [11], [60], [31], and [14], is not publicly available, preventing direct comparisons on the same dataset. We built three two-stage co-speech gesture videos generation baselines by pairing gesture generators with video synthesizers: EDGE[47] + Animate-anyone[20], EDGE[47] + Hallo[54], and DiffGest[63] + MimicMo[60]. EDGE[47] adopts the S2G-MDD[17] strategy, using DWPose to extract keypoints from the PATS[3, 2, 4] corpus and pairing them with audio for training. EDGE + Hallo further incorporates Animate-anyone’s[20] pose-driven guider and replaces Hallo’s[54] face-only static masks with region-specific upper-body masks produced by YOLOv5[40].
Co-speech Gesture Videos Generation. As summarized in Table II, our MMGT establishes a well-balanced new state of the art. It reduces video-level FVD to 230.89, 6 % lower than the previous best (DiffGest + MimicMo) and nearly half of EDGE + Hallo, while also delivering the highest PSNR and the lowest LPIPS. Its SSIM of 0.6378 is statistically tied for the lead. Lip-sync accuracy tops both measures (Syn-C, Syn-D). Facial quality is unmatched across all face metrics. Gesture realism is reflected in the best FGD. Although the Div metric may not perfectly reflect pose generation quality when significant jitter is present, it is less relevant in this context. This is further supported by the qualitative experiments in Fig. 8.
Pose Video-driven Generation. During inference, we replaced the audio-generated pose videos with full-body pose videos extracted from the PATS corpus using DWPose[56]. As shown in Table III, our model (MMGT) delivers the strongest overall performance compared with Animate-anyone [20] and MimicMo [60]. It achieves the top scores on nearly all facial metrics (FID, FVD, PSNR, LPIPS) with a competitive SSIM, while also leading the video-level FID, PSNR, and LPIPS, and noticeably reducing FVD relative to Animate-anyone. Qualitative comparisons in Fig. 9 further demonstrate that MMGT maintains body structure and identity fidelity.
@@
TABLE V: The quantitative results of the ablation study highlight the impact of the key modules on gesture generation, video quality and audiovisual synchronization metrics.
@@
Qualitative Comparative Experiment. Fig. 8 gives a qualitative comparison of o-speech gestures video generation models on the PATS dataset. The baseline pipelines (MM-Difuusion, EDGE + Animate-anyone, EDGE + Hallo, and DiffGest + MimicMo) display delayed, unnatural hand motions and noticeable lip-sync errors, along with facial distortions and background artifacts. In contrast, the videos generated by our method are highly consistent with real data, while maintaining smooth gestures and time alignment, accurate lip-syncing, and no artefacts in the background.
We also provide a qualitative comparison of Pose-driven videos generation models on the PATS dataset, as shown in Fig. 9, where Animate-anyone[20] exhibits obvious hand-shape drift and palm artefacts (yellow boxes), and its mouth region intermittently blurs, breaking identity consistency. MimicMo[60] alleviates some distortions but still shows wrist jitter and soft facial details. In contrast, the video clips generated by our model ensure that the wrist trajectory accurately follows the target posture. At the same time, facial expressions (including subtle movements of the eyes and lips) remain clear and stable over time.
@@
Lips Synchronization with Audio. For the qualitative comparison of lip synchronization, Fig. 8 shows the lip synchronization results for the word ’negative’ in the audio phrase. Our method (MMGT) generates lip shapes that closely match the Ground Truth, accurately capturing both the timing and subtle dynamics of the word. In contrast, EDGE + Animate-anyone [47, 20], DiffGest + MimicMo[63, 60], and MM-Diffusion[43] display noticeable mismatches in lip movements, leading to poor synchronization with audio. EDGE + Hallo [47, 54] exhibits significant distortions and artifacts in the lips. Additionally, we provide quantitative metrics for lip synchronization in Table II and Table V, which include Syn-C and Syn-D to evaluate synchronization accuracy and diversity, respectively. These results further confirm the robustness and superiority of our method over the baseline and ablation variants.
@@
Figure 12: Presentation and qualitative comparison of overall head diversity maps based on the PATS dataset.
@@
Effectiveness of Motion Masks in Video Generation. Fig. 11 shows a comparison of video generation results under the same input conditions using still masks and motion masks. Two key frames are selected to demonstrate the effectiveness of the proposed method. As shown, adding motion masks significantly improves the quality of generated videos, especially in the face and lip areas, where finer details and movement synchronization are better captured. By leveraging the dynamic information from motion masks, the model produces more realistic and coherent videos compared to using still masks. To give a thorough evaluation, we also confirm these findings with ablation experiments using quantitative metrics, as shown in Table V.
@@
Evaluation of Facial Motion Diversity and Synchronization. We performed a detailed comparison of facial expressions motion diversity and synchronization across individual test videos, evaluating Ground Truth, EDGE + Animate-anyone [47, 20], EDGE + Hallo [47, 54], our method MMGT, and an ablation version of our method (MMGT w/o SMGA and MMGT w/o $\mathcal{L}_{f}$). As shown in Fig. 12, the motion diversity heatmap (DIV.IMG) on the left visualizes the magnitude of the motions, with blue indicating the smallest motions, green and yellow showing moderate motions, and red highlighting the most significant motions. From the comparison of the diversity maps, our method detects the most prominent and accurate lip movements that closely match the ground truth.
@@
Detailed Comparison of Hand Motion Across Methods. We highlight the performance of different methods in reproducing hand motion. As shown in columns 2, 5, and 8 of Fig. 8, MM-Diffusion [43] generates noticeable artifacts and blurred hand shapes, failing to capture the fine details and consistency of hand movements. Similarly, EDGE + Animate-anyone, EDGE + Hallo, and DiffGest + MimicMo struggle with accurate hand positioning and shape generation, resulting in significant differences from the Ground Truth. In contrast, our method demonstrates superior fidelity by accurately capturing realistic hand gestures with precise motion and clear structural details, closely aligning with the Ground Truth.
Figure 13: We test the generalisability of our model using languages such as French and Afrikaans, respectively.
Cross-lingual Generalization. If only French or Afrikaans speech is used to replace the audio track, and not paired video ground truths, MMGT still generates lip shapes and rhythmic hand gestures that match the prosody. Notably, changes in speaking rate are naturally reflected in the frequency of mouth closures and the duration of gesture pauses, suggesting that our model operates in a language-independent manner and is highly generalisable. Due to the lack of real videos in these languages, we currently present qualitative results, as shown in Fig. 13, and plan to evaluate them manually in future work.
Generalisation to Unseen Speakers and Multiple Scenes. We trained MMGT using the TED-72h video dataset and evaluated it on unseen speakers. As shown in Fig. 14, the three rows correspond to indoor natural light, stage lighting, and a kitchen scene with deep depth of field, respectively. Under these heterogeneous backgrounds and camera positions, MMGT is able to preserve clothing textures, skin tones, and gestures; images remain stable, and facial details are consistent, indicating that the technique can be directly applied to real-world scenes and unseen speakers.
@@
Table V reports a comprehensive ablation in which we disable, one at a time, the major architectural and training ingredients of MMGT. Metrics cover gesture realism (FGD ↓, Div. ↑), facial quality, lip–audio alignment, and full-video fidelity.
Spatial Mask-guided Two-branch structure (SMT). Removing SMT (“w/o SMT”) causes the sharpest overall degradation. FGD rises from 6.02→6.78 and facial FVD more than doubles (99.38→262.20), while Syn-C/Syn-D indicate poorer lip timing (11.83 ↑ / 2.25 ↓). These drops confirm SMT’s role in coordinating coarse body motion with fine facial cues. Gesture and facial quality also decreased, as evidenced by higher FGD and FVD scores, highlighting the importance of SMGA in maintaining robust lip synchronization (Fig. 11 and Fig. 12).
Feature-wise Linear Modulation (FiLM). Excluding FiLM (“w/o FiLM”) moderately harms most face–video measures (e.g., facial FVD 99.38→132.56, video FVD 230.89→238.26), even though gesture diversity Div. improves slightly. The result suggests FiLM chiefly refines appearance rather than motion amplitude.
Figure 14: MMGT(Ours) supports many different scenarios and unseen speakers.
@@
Facial Motion Loss $\mathcal{L}_{f}$. Similarly, substituting the motion-specific facial loss $\mathcal{L}_{f}$ with a generalized loss leads to a pronounced decrease in lip synchronization accuracy. The Syn-C metric rises from 10.16 to 11.11, and the Syn-D metric increases from 4.51 to 4.65, signifying diminished articulation precision and temporal coherence between the lips and audio. As shown in Fig. 12, this degradation visibly manifests as weaker lip movements aligned with the spoken content.
Motion Mask in MM-HAA. Replacing the dynamic mask with none (“w/o Motion Mask”) boosts Syn-D to the best value (4.65), but facial and video FVDs climb sharply (205.08 and 378.32), showing that the mask primarily stabilizes global coherence and texture.
@@
Audio Guidance in Stage II. Removing the audio input entirely and relying solely on keypoint-driven generation leads to substantial performance drops in both facial and overall metrics (e.g., FVD and FID), emphasizing the essential role of audio in enhancing visual quality and temporal consistency. Furthermore, the notable decrease in lip synchronization highlights the necessity of audio-based guidance for achieving realistic, accurate lip movements.
@@
Replace Motion Mask by Still Mask. When a still mask replaces the motion mask in the MM-HAA module, face-specific metrics such as SSIM and PSNR decrease (from 0.3945 to 0.3775 and from 29.47 to 29.42, respectively), while the facial FVD score increases from 99.38 to 267.10, indicating a marked loss of detail in the face region. Gesture coherence also degrades (FGD rises from 7.04 to 7.089), and the overall video FVD increases from 230.89 to 286.389, pointing to diminished temporal consistency and video quality (Fig. 10).
With all components active, MMGT attains the statistically tied best scores on 13 of 16 indicators (e.g., facial FID 3.43, video PSNR 31.29 dB, Syn-C 10.16), validating that each module contributes to natural, temporally consistent, and visually faithful co-speech gesture video synthesis.